# Cascade Speculative Drafting for Even Faster LLM Inference

### 18 Dec 2023 v1

@@ 3.4 Algorithm
Combining the horizontal and vertical cascades, the algorithm is shown in Algorithm 1. The center of it has the a for loop which performs the horizontal cascade while the recursive call forms the vertical cascade. Noticeably, the algorithm with the last element of the draft models being the MaG algorithm, there will be no neural language models generate autoregressively.
@@ 4 Theoretical Analysis
We perform theoretical analysis to provide simple heuristics for evaluating a CS. Drafting system, and to support the generality of CS. Drafting, we prove the optimality of horizontal cascade.
@@ 4.1 Expected Walltime Improvement
We provide the formula for EWIF in an extended setting when multiple draft models are involved.
@@
the expected walltime improvement factor(EWIF) by Algorithm 1 is $T(k,{\alpha_{1},...,\alpha_{k}},{c_{1},...,c_{k}})=\frac{\sum_{i=0}^{k}\prod_{j=1}^{i}\alpha_{j}}{1+\sum_{i=1}^{k}c_{i}}$
@@
Since every token will only be considered by target model if all tokens before are accepted, a simple recursion leads to the probability of acceptance of the $i$-th token being $\prod_{j=1}^{i-1}\alpha_{j}$, and the $\prod_{j=1}^{0}\alpha_{j}=1$ representing the token generated by the target model. In the formula of the EWIF, the numerator is the expected number of tokens generated. The denominator is the time costs for each step of such a system. The formula is then the EWIF. ∎
@@
When $q_{1}=q_{2}=...=q_{k}$, let $\alpha_{i}=\alpha$ and $c_{i}=c$ for any ${}_{1}$, the EWIF is $\frac{1-\alpha^{k+1}}{(1-\alpha)(ck+1)}$
@@
The probability of $i$-th token being accepted is $\prod_{j=1}^{i}\alpha_{j}$. The derivative of EWIF with respect to $\alpha_{l}$ is $\frac{dT(k,{\alpha_{1},...,\alpha_{k}},{c_{1},...,c_{k}})}{d\alpha_{l}}=\frac{\sum_{i=l}^{k}\prod_{j=1,j\neq l}^{i}\alpha_{j}}{1+\sum_{i=1}^{k}c_{i}}$. Specifically, $\frac{dT(k,{\alpha_{1},...,\alpha_{k}},{c_{1},...,c_{k}})}{d\alpha_{1}}=\frac{\sum_{i=1}^{k}\prod_{j=2}^{i}\alpha_{j}}{1+\sum_{i=1}^{k}c_{i}}$ and $\frac{dT(k,{\alpha_{1},...,\alpha_{k}},{c_{1},...,c_{k}})}{d\alpha_{k}}=\frac{\prod_{j=1}^{k-1}\alpha_{j}}{1+\sum_{i=1}^{k}c_{i}}$
Using the information provided by (Leviathan et al., 2023), we calculated a simulated EWIF with the assumption that the event of acceptance by the target model is a Bernoulli trial. Results are shown in table 1 which shows that the cascade speculative sampling achieved better EWIF over vanilla speculative under the assumption.
4.2 Generality of Horizontal Cascade
The setting of cascade drafting can be simply extended using model interpolation that involves choosing between two models using randomness. Given two draft models $d_{1}$ and $d_{2}$ and a probability function $f$, $q_{i}$ is considered to be $f{i}d_{1}+(1-f(i))d_{2}$i.e with probability $f(i)$, $q_{i}$ is $d_{1}$, with probability $1-f{i}$ $q_{i}$ is $d_{2}$, as a result, $\alpha_{i}=f(i)\alpha_{d1}+(1-f(i))\alpha_{d2}=\alpha_{d2}+f(i)(\alpha_{d1}-\alpha_{d2})$ and $c_{i}=f(i)c_{d1}+(1-f(i))c_{d2}=c_{2}+f(i)(c_{d1}-c_{d2})$. In theorem 4.4, we prove that the optimal can always be reached without interpolation.
@@ 4.2 Generality of Horizontal Cascade
Given an cascade drafting system with the setting mentioned above. There exists a probability function $f$ that optimizes the system such that $f$ satisfies $f(i)\in\{0,1\}$ for any $i$. This means each of the tokens will be generated by either $d_{1}$ or $d_{2}$, and the optimal system is deterministic.
Proof of Theorem 4.4 is included in appendix B
Dataset $\mathcal{M}_{d1}$ $\mathcal{M}_{d2}$ $k$ $k_{2}$ EWIF
CNNDM T5-small - 9 - 2.65
CNNDM T5-base - 8 - 2.96
CNNDM T5-base T5-small 5 3 3.03
ENDE T5-small - 12 - 3.61
ENDE T5-base - 11 - 3.75
ENDE T5-base T5-small 5 8 3.93
Table 1: Simulated EWIF under the assumption that the acceptance distribution is a Bernoulli distribution. In the simulation, cascade drafting exceeded the performance of the speculative decoding on both CNN Dailymail and WMT EnDe datasets.
@@ 5 Evaluation
In addition, a recent analysis suggests GPU speed can vary for the same GPU model with one being 1.5x faster than another GPU of the same model (Sinha et al., 2022), and even for the same GPU, the performance variation between different runs is unknown; therefore, theoretical walltime improvement up to 1.5x can be undermined by GPU variation, and methods without any improvement can be measured with an improvement up to 1.5x by running on a different GPU. The properties of such lack of standardization and stability make walltime unfit for benchmarking algorithmic improvement. To resolve this issue, we designed a new benchmarking method to measure speculative decoding and its related algorithms which can reduce the effect of different/unstable GPU performance.
Our proposed method, standardized walltime improvement(SWI) calculates the GPU times of the models, assuming that each run of a language model costs the same amount of time, an assumption made when inventing the speculative sampling algorithm (Chen et al., 2023). Given a speculative decoding system $S$ containing k draft models $\mathcal{M}_{d1},...,\mathcal{M}_{dk}$, target model $\mathcal{M}_{t}$ solving a task $T$, and let $count(S,m,T)$ be the count of runs of model $m$ during the execution of task $T$ with system $S$, let $c_{i}$ be the expected cost coefficient being the time ratio between a single run of $\mathcal{M}_{i}$ and $\mathcal{M}_{t}$, and let $n$ be the total of tokens generated, the standardized walltime improvement is defined as
$\begin{split}SWI&(\mathcal{M}_{d1},...,\mathcal{M}_{dk},\mathcal{M}_{t},T)=\\ &\frac{n}{count(S,\mathcal{M}_{t},T)+\sum_{i=1}^{n}count(S,\mathcal{M}_{di},T)c_{i}}\end{split}$
which is simply the sum of GPU time by all models when $c_{i}$ is measured during a local experiment. The benefit of the $SWI$ evaluation is that $count(S,\mathcal{M}_{t},T)$ and $count(S,\mathcal{M}_{di},T)$ will be the same across different systems, so by using the same $c_{i}$ across different experiments the GPU performance variation will be eliminated; therefore, it provides a standardized, reproducible benchmark.
During our experiment, we will use two sets of $\{c_{i}\}$, existing $\{c_{i}\}$ reported by (Leviathan et al., 2023) and an heuristic $\{c_{i}\}$ being the ratio of number of tunable parameters between $\mathcal{M}_{di}$ and $\mathcal{M}_{t}$.
@@
GSM8K C.S. Drafting base, small, mag 3.75 7 1 3 1.5
@@
MMLU C.S. Drafting base, small, mag 4.83 5 1 14 2
@@
GSM8K C.S. Drafting base, small, mag 3.29 5 1 4 1.5
@@
MMLU C.S. Drafting base, small, mag 4.23 5 1 8 1.5
Table 2: The experiment results from dataset GSM8K and MMLU. BASE and SMALL refer to FLAN-T5-base and FLAN-T5-small. SWI(model size) is the standardized walltime improvement with the assumption the latency of each run of a model is its number of parameters(MS) and time cost data from previous work(PW)(Leviathan et al., 2023) respectively.
@@ 6.1 Setup
We choose our target and draft models from FLAN-T5 (Chung et al., 2022) family for our experiment as there are large variations in model sizes in the FLAN-T5 family (from 77 million to 11 billion); also, the reasoning ability of the FLAN-T5 models providing more options for our experiments. To make our experiment similar to real-world usage, We do not perform fine-tuning and the generation is done in a zero-shot manner. We use FLAN-T5-xxl as our target model, and T5-base and T5-small as our reviewing draft models and Max-Gram algorithm to be the generating draft model.
With limited computational budge, we are only able to perform sampling in temperate 00 (greedy decoding) due to the extra steps of hyperparameter tuning needed for the Cascade Speculative Decoding algorithm; the reason behind this is the benefit of temperate 00 that we can reuse the target model output which is deterministic which saves a great amount of computational resource on running the target models repeatedly.
Cascade Speculative Drafting Inference To simplify the number of hyperparameters to tune, we use MaG to generate 10 tokens at once as it’s rare that more than 10 tokens will be accepted, and we do not use lenience when the reviewer is $\mathcal{M}_{t}$ to ensure the output distribution remains do not change, and we also eliminate lenience between $\mathcal{M}_{d2}$ and MaG due to the performance map between them are large.
With these constraints, we are left with at most four parameters $k_{11}$, $k_{21}$, $k_{12}$, $l$. For the step of CS Drafting when the target model is the reviewer $k_{11}$ and $k_{12}$ are used. $k_{21}$, $l$ in the step when $\mathcal{M}_{d1}$ is the reviewer. However, there are still too many hyperparameters for us to search for the optimal result exhaustively; instead, we use halving grid search to search for a set of hyperparameters to compare to the baseline.
Baseline We consider speculative decoding and vanilla autoregressive generation as our baseline. For speculative decoding, we consider T5-small and T5-base as the draft model as they perform better than T5-large in previous work(Leviathan et al., 2023) and our preliminary experiment. We try a large range of k-values from 2 to 30 to ensure that we will find the best result for each model as the draft model to serve as our baseline.
Dataset We chose reasoning tasks for our experiment as they present more challenges of real-world NLP applications than summarization tasks. For both datasets, we experiment in a zero-shot chain-of-thought (Wei et al., 2023) setup.
The datasets we experimented on are the following:
GSM8k (Cobbe et al., 2021) A challenging mathematical reasoning dataset many large language models struggle with.
MMLU (Hendrycks et al., 2021) A reasoning dataset in many different subjects. MMLU does not have training sets meaning applying task-specific methods such as knowledge distillation to improve speculative decoding can be challenging.
@@ 6.2 Experimental Results
The experiments are shown in table 2. In two settings of SWI and Walltime, Cascade Speculative Decoding has exceeded the speculative decoding algorithm. For GSM8K, Cascade Speculative Decoding has a maximum improvement of $37\%$ over the fastest speculative algorithm; for MMLU, the maximum improvement of Cascade Speculative Decoding over speculative decoding is $72\%$.
One other observation is that despite during speculative decoding t5-small has out-performed t5-base on some experiments, but in CS Drafting with only one neural model and MaG, T5-base consistently outperforms t5-small.
@@ 7 Related Work
7.1 Efficienct Methods for Natural Language Inference
@@ 8 Conclusion
In our work, we proposed a novel algorithm CS. Drafting that involves two cascades, the vertical cascade, and the horizontal cascade. The vertical cascade eliminates the necessity of autoregressive generation from a neural language model, and the horizontal cascade effectively allocates the cost for drafting tokens at different positions. Our theories proved the optimality of the horizontal cascade, and empirically, we showed that CS. Drafting achieved up to 72 percent more speedup over speculative decoding without any additional training while maintaining generation quality.

### 21 Dec 2023 v2

@@ 3.4 Algorithm
Combining the horizontal and vertical cascades, the algorithm is presented in Algorithm 1. At its center, there is a for-loop that performs the horizontal cascade, while the recursive call forms the vertical cascade. Notably, the last element of the draft models is the MaG model.
The algorithm requires a two-dimensional hyperparameter, $K_{nn}$, with each row serving as the stop criteria for a layer of recursive calls. For simplicity, we assume the leniency parameter $l$ is universal for the algorithm, except when the target model is under review; thus, the algorithm can benefit from the speedup of leniency without altering the output distribution.
@@ 4 Theoretical Analysis
In this section, we present some theoretical analyses of the proposed Cascade Speculative Drafting algorithm.
@@ 4.1 Expected Walltime Improvement
To assist our theoretical analysts, we establish the notions. Let $\mathcal{M}_{t}$ be the target model, $\{\mathcal{M}_{i}\}$ be the draft models assisting generation with $\mathcal{M}_{i}$ being the draft model generating the $i$-th token. In the simpler case of the speculative decoding, $\mathcal{M}_{i}=\mathcal{M}_{d}$ for any $i$. Let $x$ be the input to the model at a single run, $\mathcal{M}_{t}(x)$ and $\mathcal{M}_{i}(x)$ are then the output probability distribution with input $x$.
Cost coefficient $c_{i}$ is the ratio of time for a single run of draft model $\mathcal{M}_{i}$ over target model $\mathcal{M}_{t}$.
Expected walltime improvement factor (EWIF) is the expected time improvement of an algorithm based on. We provide the formula for EWIF in an extended setting when multiple draft models are involved.
@@
The expected walltime improvement factor (EWIF) by Algorithm 1 is $T(k,{\alpha_{1},...,\alpha_{k}},{c_{1},...,c_{k}})=\frac{\sum_{i=0}^{k}\prod_{j=1}^{i}\alpha_{j}}{1+\sum_{i=1}^{k}c_{i}}$.
@@
Since every token will only be considered by the target model if all tokens before are accepted, a simple recursion leads to the probability of acceptance of the $i$-th token being $\prod_{j=1}^{i-1}\alpha_{j}$, and the $\prod_{j=1}^{0}\alpha_{j}=1$ representing the token generated by the target model. In the right-hand side, the numerator is the expected number of tokens generated, and the denominator is the time costs for each step of such a system, so the right-hand side is the expected efficiency improvement. ∎
@@
When $q_{1}=q_{2}=...=q_{k}$, let $\alpha_{i}=\alpha$ and $c_{i}=c$ for any $i$, the EWIF is $\frac{1-\alpha^{k+1}}{(1-\alpha)(c\cdot k+1)}$.
@@
The probability of $i$-th token being accepted is $\prod_{j=1}^{i}\alpha_{j}$. The derivative of EWIF with respect to $\alpha_{l}$ is $\frac{dT(k,{\alpha_{1},...,\alpha_{k}},{c_{1},...,c_{k}})}{d\alpha_{l}}=\frac{\sum_{i=l}^{k}\prod_{j=1,j\neq l}^{i}\alpha_{j}}{1+\sum_{i=1}^{k}c_{i}}$. Specifically, $\frac{dT(k,{\alpha_{1},...,\alpha_{k}},{c_{1},...,c_{k}})}{d\alpha_{1}}=\frac{\sum_{i=1}^{k}\prod_{j=2}^{i}\alpha_{j}}{1+\sum_{i=1}^{k}c_{i}}$ and $\frac{dT(k,{\alpha_{1},...,\alpha_{k}},{c_{1},...,c_{k}})}{d\alpha_{k}}=\frac{\prod_{j=1}^{k-1}\alpha_{j}}{1+\sum_{i=1}^{k}c_{i}}$.
Using the information provided by Leviathan et al. (2023), we calculat a simulated EWIF under the assumption that the event of acceptance by the target model is a Bernoulli trial. The results, shown in Table 1, indicate that speculative sampling with a horizontal cascade achieved better EWIF than vanilla speculative sampling under this assumption.
4.2 Analysis of Model Interpolation
The setting of cascade drafting can be simply extended using model interpolation that involves choosing between two models using randomness. Given two draft systems $d_{1}$ and $d_{2}$ (They can be two models generating autoregressive or two speculative decoding systems as used in vertical cascade) and a probability function $f$, an interpolation $d_{inter}$ is defined as $f(i)d_{1}+(1-f(i))d_{2}$. In other words, with probability $f(i)$, $d_{inter}$ is $d_{1}$, and with probability $1-f(i)$, $d_{inter}$ is $d_{2}$. As a result, $\alpha_{i}=f(i)\alpha_{d1}+(1-f(i))\alpha_{d2}=\alpha_{d2}+f(i)(\alpha_{d1}-\alpha_{d2})$ and $c_{i}=f(i)c_{d1}+(1-f(i))c_{d2}=c_{2}+f(i)(c_{d1}-c_{d2})$. An interpolation system can provide a smooth transition between two systems; however, in theorem 4.4, we prove that the optimal can always be reached without interpolation.
@@ 4.2 Analysis of Model Interpolation
Given a horizontal cascade system with the setting mentioned above. There exists a probability function $f$ that optimizes the system such that $f$ satisfies $f(i)\in\{0,1\}$ for any $i$. This means each of the tokens will be generated by either $d_{1}$ or $d_{2}$, and the optimal system is deterministic.
Proof of Theorem 4.4 is included in appendix B.
Dataset $\mathcal{M}_{d1}$ $\mathcal{M}_{d2}$ $k_{1}$ $k_{2}$ EWIF
CNNDM small - 9 - 2.65
CNNDM base - 8 - 2.96
CNNDM base small 5 3 3.03
ENDE small - 12 - 3.61
ENDE base - 11 - 3.75
ENDE base small 5 8 3.93
Table 1: Simulated EWIF under the assumption that the acceptance distribution is a Bernoulli distribution. BASE and SMALL refer to FLAN-T5-base and FLAN-T5-small. In the simulation, speculative sampling with horizontal cascade exceeded the performance of the vanilla speculative decoding on both CNN Dailymail (Nallapati et al., 2016) and WMT EnDe (Bojar et al., 2018) datasets.
@@ 5 Evaluation
In addition, a recent analysis suggests GPU speed can vary for the same GPU model with one being 1.5x faster than another GPU of the same model (Sinha et al., 2022), and even for the same GPU, the performance variation between different runs is unknown; therefore, theoretical walltime improvement up to 1.5x can be undermined by GPU variation, and methods without any improvement can be measured with an improvement up to 1.5x by running on a different GPU. The properties of such lack of standardization and stability make walltime unfit for measuring algorithmic improvement. To resolve this issue, we design a new metric to measure speculative decoding and its related algorithms which can reduce the impact of varying/unstable GPU performance.
Our proposed method, standardized walltime improvement (SWI), calculates the GPU times of the models, assuming that each run of a language model costs the same amount of time, an assumption made when inventing the speculative sampling algorithm (Chen et al., 2023). Given a speculative decoding system $S$ containing $k$ draft models $\mathcal{M}_{d1},...,\mathcal{M}_{dk}$, target model $\mathcal{M}_{t}$ solving a task $T$, and let $count(S,m,T)$ be the count of runs of model $m$ during the execution of task $T$ with system $S$, let $c_{i}$ be the expected cost coefficient being the time ratio between a single run of $\mathcal{M}_{i}$ and $\mathcal{M}_{t}$, and let $n$ be the total of tokens generated, the standardized walltime improvement is defined as
$\begin{split}SWI&(\mathcal{M}_{d1},...,\mathcal{M}_{dk},\mathcal{M}_{t},T)=\\ &\frac{n}{count(S,\mathcal{M}_{t},T)+\sum_{i=1}^{n}count(S,\mathcal{M}_{di},T)c_{i}},\end{split}$
which is simply the sum of GPU time by all models when $c_{i}$ is measured during a local experiment. The benefit of the $SWI$ evaluation is that $count(S,\mathcal{M}_{t},T)$ and $count(S,\mathcal{M}_{di},T)$ will be the same across different systems, so by using the same $c_{i}$ across different experiments the GPU performance variation will be eliminated; therefore, it provides a standardized, reproducible measurement.
During our experiment, we will use two sets of $\{c_{i}\}$, existing $\{c_{i}\}$ reported by Leviathan et al. (2023) and an heuristic $\{c_{i}\}$ being the ratio of number of tunable parameters between $\mathcal{M}_{di}$ and $\mathcal{M}_{t}$.
@@
GSM8K C.S. Drafting base, small, mag 3.75 7 1 10 1.5
@@
MMLU C.S. Drafting base, small, mag 4.83 5 1 19 2
@@
GSM8K C.S. Drafting base, small, mag 3.29 5 1 9 1.5
@@
MMLU C.S. Drafting base, small, mag 4.23 5 1 13 1.5
Table 2: The experiment results from datasets GSM8K and MMLU. BASE and SMALL refer to FLAN-T5-base and FLAN-T5-small. SWI (MS) is the standardized walltime improvement with the assumption that the latency of each run of a model is its number of parameters (model size). SWI (PW) is the SWI with the assumption that the latency of each run of a model is the time cost data reported from previous work (Leviathan et al., 2023). $k_{11}$, $k_{21}$, $k_{12}$, $l$ are the hyperparameters. $k_{1i}$ represents the step limitation target model and the draft models, $k_{21}$ is the step limitations between the first and second draft model, and $l$ is leniency as shown in algorithm 1. For speculative decoding, the $k_{11}$ is simply the $k$.
@@ 6.1 Setup
We chose our target and draft models from the FLAN-T5 (Chung et al., 2022) family for our experiment, as there is a large variation in model sizes within the FLAN-T5 family (ranging from 77 million to 11 billion parameters). To align our experiment with current common usage, we do not perform fine-tuning, and the generation is conducted in a zero-shot manner. We use FLAN-T5-xxl as our target model, FLAN-T5-base and FLAN-T5-small as our reviewing draft models, and the Max-Gram algorithm as the generating draft model.
To ensure our experiments are fully reproducible, we perform sampling at temperature 00, which is essentially greedy decoding. And we do not observe any significant difference between sampling with temperature $1$ and greedy decoding in previous speculative decoding experiments (Leviathan et al., 2023).
Cascade Speculative Drafting Inference To reduce the number of hyperparameters to tune, we use MaG to generate 10 tokens at once, as it is rare for more than 10 tokens to be accepted. We do not use lenience when the reviewer is $\mathcal{M}_{t}$ to ensure the output distribution remains does not change. Additionally, we eliminate leniency between $\mathcal{M}_{d2}$ and MaG because the performance gap between them is large.
With these constraints, we are left with at most four hyperparameters: $k_{11}$, $k_{21}$, $k_{12}$, and $l$. For the CS Drafting step where the target model is the reviewer, $k_{11}$ and $k_{12}$ are used. $k_{21}$ and $l$ are used in the step where $\mathcal{M}_{d1}$ is the reviewer. However, there are still too many hyperparameters for us to exhaustively search for the optimal result; instead, we employ a halving grid search to find a set of hyperparameters.
Baseline We consider both speculative decoding and vanilla autoregressive generation as our baselines. For speculative decoding, we opt for FLAN-T5-small and FLAN-T5-base as the draft models, as they perform better than FLAN-T5-large, according to Leviathan et al. (2023) and our preliminary experiments. We test a wide range of $k$-values, from 2 to 30, to ensure that we find the best result for each model serving as our draft model to establish our baseline.
Dataset We chose two commonly used datasets for our experiments. For both datasets, we conducted experiments in a zero-shot chain-of-thought setup (Kojima et al., 2022; Wei et al., 2023):
  • GSM8K (Cobbe et al., 2021) is a dataset comprising 8,500 high-quality, linguistically diverse, grade-school math word problems. It focuses on multi-step reasoning with problems that are typically solvable using basic arithmetic in 2-8 steps.
  • MMLU (Hendrycks et al., 2021), or Massive Multitask Language Understanding, is a benchmark for testing how well large language models grasp knowledge. It encompasses 57 diverse subjects, ranging from elementary science to advanced law.
@@ 6.2 Experimental Results
Table 2 presents the experimental results. In two settings of SWI, Cascade Speculative Decoding has outperformed the speculative decoding algorithm. For GSM8K, Cascade Speculative Decoding achieved a maximum additional speedup of 37% over the fastest speculative algorithm; for MMLU, the maximum additional speedup improvement over speculative decoding is 72%.
Additionally, we emphasize the effectiveness of the MaG algorithm. When comparing CS Drafting with one neural model and MaG against the fastest speculative decoding setup, we found that CS Drafting with one neural model gained up to a 56% speedup on MMLU and a 24% speedup on GSM8K. Notably, the MaG algorithm only involves a bigram model with as many parameters as the size of the tokenizer, making its memory cost negligible. Thus, the speedup gained using CS Drafting with one neural model involves no additional deployment overhead while reducing both latency and computational cost, making it a superior choice over speculative decoding.
When comparing CS Drafting with two neural models to CS Drafting with one neural model, the former requires additional memory utilization. Despite the latency reduction, CS Drafting with two neural models can potentially incur more computational cost. However, during our experiments, since the draft models had only 2% of the number of parameters of the target model, the memory cost did not outweigh the reduced latency. Therefore, the CS Drafting algorithm with two neural models still uses less computational resources, as it completes each generation round faster.
Another observation is that despite FLAN-T5-small sometimes outperforming FLAN-T5-base as a draft model for speculative decoding, in CS Drafting with the aid of MaG, FLAN-T5-base consistently outperforms FLAN-T5-small. This raises a topic for future study: determining the ideal size of draft models for CS Drafting.
@@ 7 Related Work
7.1 Efficienct Methods for Language Model Inference
@@ 8 Conclusion
In our work, we propose a novel algorithm, CS Drafting, which involves two cascades: the vertical cascade and the horizontal cascade. The vertical cascade eliminates the necessity of autoregressive generation from a neural language model, while the horizontal cascade effectively allocates the cost of drafting tokens at different positions. Our experiments demonstrate that CS Drafting achieves up to a 72 percent additional speedup over speculative decoding without any additional training, while maintaining the same generation quality as the target model.